From nch@roe.ac.uk Mon Jun 18 16:26:35 2007 Date: Mon, 18 Jun 2007 14:38:08 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , John Taylor , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting, 8 June 2007 Minutes of WFAU VDFS Science Archive meeting: 8th June 2007 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, JB, LGR, NJC, PMW, ETWS, MAR Apologies: JDT, BCW, JPE, AL, RGM, JMS, MSH ... and apologies for the lateness of distribution of these minutes ... DONM: 10am, Friday 22nd June 2007 in the Plate Library Actions discharged this week: ----------------------------- ACTION: NCH to discuss publicity image gallery with MAR. Discharged; MAR will look into this and discuss with various people. Actions partly discharged but continuing: ----------------------------------------- The following from last time partly done but continue: o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH Continues; will get sorted (honestly!) as part of the general ingest DB schema revamp to facilitate rapid generalised photometric/astrometric recalibration. Actions carried forward from 01/06/07 meeting: ---------------------------------------------- The following from last time continue: ACTION: NCH to discuss and review AstroWISE interfacing with MAR and JDT. ACTION: NCH to start some hard-nosed negotiations with the UKIDSS PI and CSS concerning the contents and timing of DR3. - CONTINUES; remaining problems with ZY NDR data from early 06B should be sorted quite soon now, so the schedule should become easier to tie down. ACTION: AL to check on the exact wording of the survey release policy with reference to world release of proprietary survey data. - CONTINUES Specific points and new actions: -------------------------------- Project management: NCH noted that the Rolling Grant went in on Tuesday, complete with cases for resources for operations and a final development push. Thanks are due especially to PMW and RGM for tenacious efforts with STFC JeS, UoE HR and the whole FECing procedure. PMW has prepared the progress review materials for this afternoon's VDMT and will attend (NCH is skiving). WFCAM & VISTA updates: Nothing new to report. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: JB noted: "Despite work being done on the UKLight link we've been free of downtime since it was last fixed. Transfer are up-to-date but for a few days of 06B that CASU seem to be finishing off but have yet to mark as ok to copy (excepting the first months of 06B of course). Work appears to have started at CASU on 07A however no more than the two we already have have been marked okay to copy." WSA Operations: JB reported: "Backups and CUs were run as normal. CU3 has finished almost all but the first chunk of 06B. CU2 has gotten a good way into the data we have. Ingestion has been done on a daily basis so a large tranche of 06B jpegs are now available." Hardware: MAR and NCH noted a user-reported problem with the batch server hosting the SDSS databases; JB noted: "We have what appears to be a backplane problem with Thumose however Ian from Eclipse is coming in on Friday afternoon to look at it (so slight downtime for Thutmose but hopefully it will be fixed). Ian also sent over a couple of replacement disks we were waiting on. Thursday afternoon I went ot a talk on filesystems and storage, some interesting discussion came about with the decision to start up a group for computer people interested in storage, this I will be keeping tabs on so as to benefit from their expertise :) (The main direct consequences are that I may have found a cheaper source of NAS boxes and that in the long term it looks like ZFS might be a good filesystem for us to use)." Software: NJC reported working on updating code with numpy and fixing associated bugs. This is still ongoing, as there are many bits of code which use numarray and pyfits. NJC has also been testing new changes to the half-light radius code and doing comparisons between galaxy photometry in an overlap of the DXS and LAS (in collaboration with UKIDSS user Anthony Smith). MAR reported: "Had a go at extending the GPS mosaic towards the galactic centre but results are not as good, presumably due to crowding/sky background problems. However one of the colour images produced showed up a problem with a few detector images where it looks like the WCS determination has gone wrong. Currently investigating what atrributes in CurrentAstrometry can be used to pick these up for flagging at QC and feedback to CASU." RSC and NJC noted that it is now time to get on with implementing the schema change required for fast recalibration. ETWS has some final investigations to do with CU4 which ideally should be done before branching the CVS; hence, branching will occur on Monday next week so that code changes can be made and tested against TestWSAnch which (mostly) has the new schema and contains the UKIDSS DXS data for test purposes. RSC reported tidying up code, writing off a few bugs and todo's prior to starting work on the recalibration branch. (trac tickets 26, 28, 50, 53 & 65) RSC reminded NCH that a couple of minor tweaks are required in the source merging procedure as a result of some user queries this week. RSC will make a note under required modifications for CU7 in the relevant trac ticket. ETWS reported fixing smaller inconsistencies in the ingest monitoring page cration script, helping NJC with a CasuSource CVS update, and installing the latest pyFITS release candidate and missing docutils for epydoc. Survey Data Release: Nothing to report this week. Non-survey Data Release: Nothing to report this week. Astrogrid deployment: Nothing new to report this week Miscellaneous: JB noted that RGM had circulated a heads-up concerning a local meeting with people at NeSC who run the Edinburgh Compute and Data Facility (ECDF - see http://www.is.ed.ac.uk/ECDF). RGM is inviting them up to give a coffee talk and then to talk with interested folks. Following on from emails from RGM, AL and Head of Group, NJC kindly volunteered to attend next week's SUPA jamboree on Wednesday and will present poster(s) advertising WFAU's activities within the Institute.